OpenStudy (anonymous):

A rectangular tank of height of 12cm and base 5cm by 7cm is filled with water to a height of 8cm. By how much will the water rise if a stone of volume 70cm is dropped into the tank?

OpenStudy (anonymous):

"How much will the water rise" means that the height will change. The area of the horizontal base will stay the same. A of base = 5 * 7 = 35 If you have a rectangular tank that has a base of 35cm², how tall does it need to be to store 70cm³ water? Simple: 70cm³/35cm² = 2cm. Well, that's the exact situation we have here. Adding the rock is equivalent to "adding" that tank full of water. It will rise 2cm.